# Apparate: Rethinking Early Exits to Tame Latency-Throughput Tensions in ML Serving

### v1

@@ Abstract
Machine learning (ML) inference platforms are tasked with balancing two competing goals: ensuring high throughput given many requests, and delivering low-latency responses to support interactive applications. Unfortunately, existing platform knobs (e.g., batch sizes) fail to ease this fundamental tension, and instead only enable users to harshly trade off one property for the other. This paper explores an alternate strategy to taming throughput-latency tradeoffs by changing the granularity at which inference is performed. We present Apparate, a system that automatically applies and manages early exits (EEs) in ML models, whereby certain inputs can exit with results at intermediate layers. To cope with the time-varying overhead and accuracy challenges that EEs bring, Apparate repurposes exits to provide continual feedback that powers several novel runtime monitoring and adaptation strategies. Apparate lowers median response latencies by 40.5-91.5% and 10.0-24.2% for diverse CV and NLP workloads, respectively, without affecting throughputs or violating tight accuracy constraints.
Apparate: Rethinking Early Exits to Tame Latency-Throughput Tensions in ML Serving
Yinwei ${}^{\star}$${}^{\mathsection}$ Rui ${}^{\star}$${}^{\mathsection}$ Anand ${}^{\dagger}$ Kai ${}^{\mathsection}$ Ravi ${}^{\mathsection}$
${}^{\mathsection}$rinceton University ${}^{\dagger}$eorgia Institute of Technology
00footnotetext: ${}^{\star}$ Equal contributions.
1 Introduction
@@
Machine Learning (ML) inference has become a staple for request handling in interactive applications such as traffic analytics, chatbots, and web services [40, 33, 26, 55, 24]. To manage these ever-popular workloads, applications typically employ serving platforms [22, 39, 5, 4, 44, 17] that ingest requests and schedule inference tasks with pre-trained models across large clusters of compute resources (typically GPUs). The overarching goals of serving platforms are to deliver sufficiently high throughput to cope with large request volumes – upwards of trillions of requests per day [37] – while respecting the service level objectives (SLOs) that applications specify for response times (often 10s to 100s of ms).
@@
Unfortunately, in balancing these goals, serving platforms face a challenging tradeoff (§2.1): requests must be batched for high resource efficiency (and thus throughput), but larger batch sizes inflate queuing delays (and thus per-request latencies). Existing platforms navigate this latency-throughput tension by factoring only tail latencies into batching decisions and selecting max batch sizes that avoid SLO violations. Yet, this trivializes the latency sensitivity of many interactive applications whose metrics of interest (e.g., user retention [20, 54], safety in autonomous vehicles [52]) are also influenced by how far below SLOs their response times fall.
@@
This paper explores the role that early exits (EEs) – an adaptation mechanism that has garnered substantial ML research interest in recent years [53, 58, 56, 64, 28, 57, 36] – can play in resolving this tension for existing serving platforms. With EEs, intermediate model layers are augmented with ramps of computation that aim to predict final model responses. Ramp predictions with sufficiently high confidence (subject to a threshold) exit the model, foregoing downstream layers and bringing corresponding savings in both compute and latency. The intuition is that models are often overparameterized (especially given recent model growth [48, 31, 32]), and certain ‘easy’ inputs may not require complete model processing for accurate results. Importantly, unlike existing platform knobs (e.g., batch size) that simply walk the steep latency-throughput tradeoff curve, EEs rethink the granularity of inference on a per-input basis. This, in turn, provides a path towards lowering request latencies without harming platform throughputs. Indeed, across different CV and NLP workloads, we find that optimal use of EEs brings 24.8-94.0% improvement in median latencies for the same accuracy and throughput.
@@
Despite these potential benefits, EEs are plagued with practical challenges that have limited their impact to date (§2.3). The primary issue is that EE proposals have solely come in the context of specific model architectures that impose fixed ramp designs and locations [57, 53]. The lack of guidance for integrating EEs into arbitrary models is limiting, especially given the ever-growing number of model offerings in the marketplace today. Worse, even existing proposals lack any policy for runtime adaptation of EE configurations, i.e., the set of active ramps and their thresholds. Such adaptation is crucial since dynamic workload characteristics govern the efficacy of each ramp in terms of exiting capabilities and added overheads (to non-exiting inputs); accordingly, failure to continually adapt configurations can result in unacceptable accuracy drops of 8.3-23.9% for our workloads. However, devising adaptation policies is difficult: the space of configurations is massive, and it is unclear how to obtain a signal for accuracy monitoring once an input exits.
@@
We present Apparate, the first system that automatically injects and manages EEs for serving with a wide range of models. Our main insight is that the above challenges are not fundamental to EEs, and instead are a byproduct of what we are trying to get out of them. Specifically, adaptation challenges largely stem from halting execution for an input upon an exit, which leaves uncertainty in the ‘correct’ response (as per the non-EE model). Instead, Apparate uses EEs only to deliver latency reductions; results for successful exits are immediately released, but all inputs continue to the end of the model. The key is in leveraging the (now) redundant computations to enable continual and efficient adaptation.
@@
Guided by this philosophy, Apparate runs directly atop existing serving platforms and begins by automatically converting registered models into EE variants. Apparate’s EE preparation strategy must strike a balance between supporting fine-grained runtime adaptation without burdening those time-sensitive algorithms with (likely) unfruitful options. To do so without developer effort, Apparate leans on guidance from the original model design, crafting ramp locations and architectures based on downstream model computations and data flow for intermediates around the model. Original model layers (and weights) are unchanged, and added ramps are trained in parallel (for efficiency), but in a manner that preserves their independence from other ramps.
@@
Once deployed by serving platforms, Apparate continually monitors EE operation in GPUs, tracking computations and latency effects of each ramp, as well as outputs of the original model (for accuracy ground truth). To tackle the massive space of configuration options, Apparate judiciously decouples tunable EE knobs: thresholds for existing ramps are frequently and quickly tuned to ensure consistently high accuracy, while costlier changes to the set of active ramps occur only periodically as a means for latency optimization. For both control loops, Apparate leverages several fundamental properties of EEs to accelerate the tuning process. For instance, the monotonic nature of accuracy drops (and increases in latency savings) for higher thresholds motivates Apparate’s greedy algorithm for threshold tuning which runs up to 3 orders of magnitude faster than grid search while sacrificing only 0-3.8% of the potential wins.
@@
We evaluated Apparate across a variety of recent CV and NLP models (ranging from compressed to large language models), multiple workloads in each domain, and several serving platforms (TensorFlow-Serving [39], Clockwork [22]). Compared to serving without EEs, Apparate improves 25th percentile and median latencies by 70.2-94.2% and 40.5-91.5% for CV, and 16.0-37.3% and 10.0-24.2% for NLP, while imposing negligible impact on platform throughput. Importantly, unlike existing EE proposals that yield accuracy dips of 8.3-23.9%, we find that Apparate’s adaptation strategies always met user-defined accuracy constraints.
2 Background and Motivation
@@
We start by overviewing existing ML serving platforms (§2.1), highlighting the challenges they face in balancing metrics that are important for system performance (i.e., throughput, resource utilization) and application interactivity, i.e., per-request latencies. We then describe the promising role that early exits can play in alleviating those tensions (§2.2), and the challenges in realizing those benefits in practice (§2.3). Results here follow the methodology from §5.1, and presented trends hold for all workloads used in the paper.
2.1 Model Serving Platforms
@@
ML models are routinely used to service requests in interactive applications such as real-time video analytics [49, 13], recommendation engines [50], or speech assistants [12]. To manage such workloads, especially at large scale, applications employ serving platforms such as ONNX runtime [5], TensorFlow-Serving [39], PyTorch Serve [9], Triton Inference Server [4], among others [60, 22, 17, 49, 44]. These platforms ingest pre-trained model(s) from applications, often in graph exchange formats like ONNX [6] and NNEF [2], and are also granted access to a pool of compute resources (potentially with ML accelerators such as GPUs) for inference.
@@
Given the latency-sensitive nature of interactive applications, requests are most often accompanied with service level objectives (SLOs) that indicate (un)acceptable response times for the service at hand. In particular, responses delivered after an SLO expires are typically discarded or yield severely degraded utility. Common SLOs are in the 10-100s of milliseconds, e.g., for live video analytics [49, 40].
@@
During operation, serving platforms queue up incoming requests that can arrive at fixed or variable rates, and continually schedule jobs across the available compute resources. An inference task may be scheduled to run on a single node in a cluster, or may be distributed across multiple nodes [60]. State-of-the-art serving frameworks include optimizations such as response caching [17], intelligent job placement to aggressively pack GPU memory resources [49], or strategies to mitigate CPU-GPU communication overheads [60].
@@
Figure 1: Throughput-latency tradeoff in model serving. Results show serving times with batch sizes of 1-16.
@@
Latency-Throughput tension. Real-world inference deployments must handle large volumes of requests [25, 37]. To support the need for high throughput, serving platforms resort to batching, whereby inputs are grouped into a single high-dimensional tensor that moves through the model in lockstep, kernel by kernel, with final per-request responses being delivered at the same time. Larger batch sizes amortize the cost of loading a kernel into GPU memory across more inputs, and enable more effective use of the parallelism that ML-focused hardware affords [17, 62].
Unfortunately, delivering the throughput necessary to support high request rates is directly at odds with per-request latencies (Figure 1). On one hand, latency for an input is minimized by scheduling inference as soon as the request arrives with batch size of 1. On the other hand, throughput is maximized by creating large batches using a queuing system which directly inflates request latencies.
The problem. In navigating this tension, the key decision that serving platforms face is when to drain queued requests and schedule them for inference. Certain platforms [22, 49, 17] take an all-or-nothing viewpoint on latency, with adherence to SLOs considered complete success, and violations viewed as complete failure. Accordingly, these platforms schedule inference jobs in a work-conserving manner and select the max batch size that limits SLO violations for queued requests. However, many interactive applications present a more nuanced latency story where sub-SLO responses are not equally useful. For instance, live video analytics results can change every frame, i.e. 33 ms (Figure 2); faster responses ensure up-to-date perception of the environment. Similarly, for speech assistance, faster responses are favored to keep conversational interactivity high [26, 59].
Figure 2: Object counts in videos vary frequently and thus need low-latency responses. Results use a random 30-fps video in our corpus and the FasterRCNN detector.
Other platforms [39, 4, 9] provide more flexibility by exposing tunable knobs to guide queue management. For instance, applications can configure $max\_batch\_size$ and $batch\_timeout\_micros$ parameters that set a cap on batch size or inter-job scheduling durations. However, as shown in Figure 3, such knobs do little to ease the throughput-latency tension, and instead expose harsh tradeoffs: across both CV and NLP workloads, tuning these knobs for median latency improvements of 17.3-39.1% brings 1.1-3.6$\times$ reductions in average batch sizes (and proportional hits on throughput).
@@
Figure 3: Tuning platform knobs lowers latencies but harms throughput. Results vary TF-Serve’s $max\_batch\_size$ from 4-16. Gray lines show min serving delay per model (batch of 1). CV uses a random video in our set; NLP uses Amazon reviews [10].
2.2 Early-Exit Models
Figure 4: Early exit networks enable early termination of inputs at intermediate layers, lowering both compute and latency.
Early (or multi) exit models [53, 57] present an alternate way to address this tension by rethinking the granularity of inference. As shown in Figure 4, the key premise is that certain ‘easy’ inputs may not require the full predictive power of a model to generate an accurate result. Instead, results for such inputs may be predictable from the values at intermediate layers. In such cases, the foregone model execution can yield proportional reductions in both per-request latencies and compute footprints. Thus, the goal with early exits (EEs) is to determine, on a per-input basis, the earliest model layer at which an accurate response can be generated.
@@ 2.2 Early-Exit Models
Figure 5: EEs can lower latencies without harming throughput. Results modulate latencies from TF-Serve with original/vanilla models (Figure 3) based on optimal exiting.
Potential benefits. To understand the effect that EEs can have on the latency-throughput tension, we considered off-the-shelf EE variants for the models used in Figure 3: BranchyNet [53] (for CV) and DeeBERT [57] (for NLP). For each model-input pair, we identified the optimal exit point defined as the earliest exit ramp that predicted the correct response for the input, i.e., the ramp which assigned the maximum probability to the correct label. We then modified the highest-throughput results in Figure 3 to account for exiting by subtracting the time saved for each exiting input, i.e., the difference in time for passing an input to the end of its optimal ramp versus passing it to the end of the model (without any ramps present). Note that these results are conservative upper bounds in that they do not reduce queuing delays or alter job scheduling. As shown in Figure 5, without changing queueing decisions, EEs can bring 35-54.7% and 17.9-26% improvements in median and 95th percentile latencies relative to running existing serving systems alone.
2.3 Challenges
Despite numerous EE proposal from the ML community [64, 36, 57, 28, 53, 58, 46], and their potential benefits for serving platforms, multiple issues complicate their use in practice, leading to low adoption rates. We describe them in turn.
C1: Latency and resource overheads. Although exiting can enable certain inputs to eschew downstream model computations, exit ramps impose two new overheads on model serving. First, to be used, ramps must also be loaded into GPU memory which is an increasingly precious resource as models grow in size [60, 48, 31] and inference shifts to resource-constrained settings [40, 23]. For instance, DeeBERT inflates overall memory requirements by 6.56% compared to BERT-base. Second, certain inputs may be too “hard” to accurately exit at an intermediate ramp. In these cases, overall serving latency mildly grows as unsuccessful exiting decisions are made, e.g., inputs that cannot exit at any ramp slow by 22.0% and 19.5% with BranchyNet and DeeBERT.

### v2

@@ Abstract.
Machine learning (ML) inference platforms are tasked with balancing two competing goals: ensuring high throughput given many requests, and delivering low-latency responses to support interactive applications. Unfortunately, existing platform knobs (e.g., batch sizes) fail to ease this fundamental tension, and instead only enable users to harshly trade off one property for the other. This paper explores an alternate strategy to taming throughput-latency tradeoffs by changing the granularity at which inference is performed. We present Apparate, a system that automatically applies and manages early exits (EEs) in ML models, whereby certain inputs can exit with results at intermediate layers. To cope with the time-varying overhead and accuracy challenges that EEs bring, Apparate repurposes exits to provide continual feedback that powers several novel runtime monitoring and adaptation strategies. Apparate lowers median response latencies by 40.5–91.5% and 10.0–24.2% for diverse CV and NLP classification workloads, and median time-per-token latencies by 22.6–77.9% for generative scenarios, without affecting throughputs or violating tight accuracy constraints.
copyright: acmlicensedjournal: JDSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8journalyear: 2024copyright: rightsretainedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695963isbn: 979-8-4007-1251-7/24/11 Equal contributions.
1. Introduction
@@
Machine Learning (ML) inference has become a staple for request handling in interactive applications such as traffic analytics, chatbots, and web services  (Padmanabhan et al., 2022; Porter, 2023; Hinton et al., 2012; Wu et al., 2019). To manage these ever-popular workloads, applications typically employ serving platforms (Gujarati et al., 2020; Olston et al., 2017; onn, 2024a; tri, 2024; Romero et al., 2021; Crankshaw et al., 2017; Kwon et al., 2023; Yu et al., 2022) that ingest requests and schedule inference tasks with pre-trained models across large clusters of compute resources (typically GPUs). The overarching goals of serving platforms are to deliver sufficiently high throughput to cope with large request volumes – upwards of billions of requests per day (Mattson et al., 2020; ONN, 2023) – while respecting the service level objectives (SLOs) that applications specify for response times (often 10–100s of ms).
@@
Unfortunately, in balancing these goals, serving platforms face a challenging tradeoff (§2.1): requests must be batched for high resource efficiency (and thus throughput), but larger batch sizes inflate queuing delays (and thus per-request latencies). Existing platforms navigate this latency-throughput tension by factoring only tail latencies into batching decisions and selecting max batch sizes that avoid SLO violations. Yet, this trivializes the latency sensitivity of many interactive applications whose metrics of interest (e.g., user retention (Gigaspaces, 2023; Think with Google, 2017), safety in autonomous vehicles (Tanwar et al., 2019)) are also influenced by how far below SLOs their response times fall.
@@
This paper explores the role that early exits (EEs) – an adaptation mechanism that has garnered substantial ML research interest in recent years (Teerapittayanon et al., 2017; Xin et al., 2021; Xie et al., 2021; Zhou et al., 2020; Huang et al., 2018a; Xin et al., 2020; Liu et al., 2020) – can play in resolving this tension. With EEs, intermediate model layers are augmented with ramps of computation that aim to predict final model responses. Ramp predictions with sufficiently high confidence (subject to a threshold) exit the model, foregoing downstream layers and bringing corresponding savings in both compute and latency. The intuition is that models are often overparameterized (especially with recent growth (Sevilla et al., 2021; Jeon et al., 2019)), and ‘easy’ inputs may not require complete model processing for accurate results. Importantly, unlike existing platform knobs (e.g., batch size) that simply walk the steep latency-throughput tradeoff curve, EEs rethink the granularity of inference on a per-input basis. This provides a path towards lowering request latencies without harming platform throughputs. Indeed, across CV and NLP workloads, we find that optimal use of EEs brings 24.8–94.0% improvement in median latencies for the same accuracy and throughput.
@@
Despite the potential benefits, EEs are plagued with practical challenges that have limited their impact (§2.3). The primary issue is that EE proposals have solely come in the context of specific model architectures that impose fixed ramp designs and locations (Xin et al., 2020; Teerapittayanon et al., 2017; Schuster et al., 2022; Bae et al., 2023). The lack of guidance for integrating EEs into arbitrary models is limiting, especially given the ever-growing model offerings in the marketplace. Worse, even existing proposals lack any policy for runtime adaptation of EE configurations, i.e., the set of active ramps and their thresholds. Such adaptation is crucial since dynamic workload characteristics govern the efficacy of each ramp in terms of exiting capabilities and added overheads (to non-exiting inputs); failure to continually adapt configurations can result in unacceptable accuracy drops up to 23.9% for our workloads. However, devising adaptation policies is difficult: the space of configurations is massive, and it is unclear how to obtain a signal for accuracy monitoring once an input exits.
@@
We present Apparate, the first system that automatically (i.e., without developer effort or expertise) injects and manages EEs for serving with a wide range of models. Our main insight is that the above challenges are not fundamental to EEs, and instead are a byproduct of what we are trying to get out of them. Specifically, adaptation challenges largely stem from halting execution for an input upon an exit, which leaves uncertainty in the ‘correct’ response (as per the non-EE model). Instead, Apparate uses EEs only to deliver latency reductions; results for successful exits are immediately released, but all inputs continue to the end of the model. The key is in leveraging the (now) redundant computations to enable continual and efficient adaptation, while also remaining compatible with proven compute efficiency optimizations such as batching and model compression.
@@
Guided by this philosophy, Apparate runs directly atop existing serving platforms and begins by automatically converting registered models into EE variants. Apparate’s EE preparation strategy must strike a balance between supporting fine-grained runtime adaptation without burdening those time-sensitive algorithms with (likely) unfruitful options. To do so without developer effort, Apparate leans on guidance from the original model design, crafting ramp locations and architectures based on downstream model computations and data flow for intermediates around the model. Original model layers (and weights) are unchanged, and added ramps are rapidly trained in parallel (for efficiency), but in a manner that preserves their independence from other ramps.
@@
Once deployed, Apparate continually monitors EE operation in GPUs, tracking computations and latency effects of each ramp, as well as outputs of the original model (for accuracy ground truth). To tackle the massive space of configuration options, Apparate judiciously decouples tunable EE knobs: thresholds for existing ramps are frequently and quickly tuned to ensure consistently high accuracy, while costlier changes to the set of active ramps occur only periodically as a means for latency optimization. For both control loops, Apparate leverages several fundamental properties of EEs to accelerate the tuning process. For instance, the monotonic nature of accuracy drops (and increases in latency savings) for higher thresholds motivates Apparate’s greedy algorithm for threshold tuning which runs 3 orders of magnitude faster than grid search while sacrificing only 0–3.8% of the potential latency wins.
@@
We evaluated Apparate across a variety of recent CV and NLP models (ranging from compressed to large language models), diverse workloads (classification and generative), and several serving platforms (TensorFlow-Serving (Olston et al., 2017), Clockwork (Gujarati et al., 2020), and HuggingFace Pipelines (hf_, 2024)). Compared to serving without EEs, Apparate improves 25th percentile and median classification latencies by 70.2–94.2% and 40.5–91.5% for CV, and 16.0–37.3% and 10.0–24.2% for NLP, while imposing negligible impact on platform throughput. Latency wins are similar for generative scenarios: 22.6–77.9% median time-per-token improvements. Importantly, unlike existing EE proposals that yield accuracy dips up to 23.9%, we find that Apparate’s adaptation strategies always met set accuracy constraints. We open source Apparate at https://github.com/dywsjtu/apparate.
2. Background and Motivation
@@
We start by overviewing existing ML serving platforms (§2.1), highlighting the challenges they face in balancing metrics that are important for system performance (i.e., throughput, resource utilization) and application interactivity, i.e., per-request latencies. We then describe the promising role that early exits can play in alleviating those tensions (§2.2), and the challenges in realizing those benefits in practice (§2.3). Results follow the methodology from §4.1.
2.1. Model Serving Platforms
@@
ML models are routinely used to service requests in interactive applications such as real-time video analytics (Shen et al., 2019; Bhardwaj et al., 2022), chatbots (Touvron et al., 2023), recommendation engines (Sima et al., 2022), or speech assistants (Ahmed et al., 2020). To manage such workloads, especially at large scale, applications employ serving platforms such as ONNX runtime (onn, 2024a), TensorFlow-Serving (Olston et al., 2017), PyTorch Serve (Tor, 2024), Triton Inference Server (tri, 2024), among others (Yu et al., 2022; Gujarati et al., 2020; Crankshaw et al., 2017; Shen et al., 2019; Romero et al., 2021; Kwon et al., 2023). These platforms ingest pre-trained model(s), often in graph exchange formats like ONNX (onn, 2024b) and NNEF (nne, 2018), and are granted access to a pool of compute resources (usually with ML accelerators such as GPUs) for inference.
@@
Given the latency-sensitive nature of interactive applications, requests are often accompanied with service level objectives (SLOs) that indicate (un)acceptable response times for the service at hand. In particular, responses delivered after an SLO expires are typically discarded or yield severely degraded utility. Common SLOs are in the 10–100s of milliseconds, e.g., for live video analytics (Shen et al., 2019; Padmanabhan et al., 2022).
@@
During operation, serving platforms queue up incoming requests that can arrive at fixed or variable rates, and continually schedule jobs across the available compute resources. An inference task may be scheduled to run on a single node in a cluster, or may be distributed across multiple nodes (Yu et al., 2022; Gujarati et al., 2020).
@@
Figure 1. Throughput-latency tradeoff in model serving. Results show serving times with batch sizes of 1–16.
@@
Latency-Throughput tension. To support the need for high throughput, serving platforms resort to batching, whereby inputs are grouped into a single high-dimensional tensor that moves through the model in lockstep, kernel by kernel, with final per-request responses being delivered at the same time. Larger batch sizes amortize the cost of loading a kernel into GPU memory across more inputs, and enable more effective use of accelerator parallelism (Crankshaw et al., 2017; Zhang et al., 2019).
@@
Figure 2. Tuning platform knobs lowers latencies but harms throughput. Results vary TF-Serve’s $max\_batch\_size$ from 4–16. Gray lines show min serving time per model (batch=1). CV uses a random corpus video; NLP uses Amazon reviews (ama, 2013).
Figure 3. Early exits enable termination of inputs at intermediate layers, lowering both compute and latency.
Unfortunately, delivering the throughput necessary to support high request rates (Mattson et al., 2020; Hazelwood et al., 2018) is directly at odds with per-request latencies (Figure 1). On one hand, latency for an input is minimized by scheduling inference as soon as the request arrives with batch size of 1. On the other hand, throughput is maximized by creating large batches using a queuing system which directly inflates request latencies.
The problem. In navigating this tension, the key decision that serving platforms face is when to drain queued requests for inference. Certain platforms (Gujarati et al., 2020; Shen et al., 2019; Crankshaw et al., 2017) take an all-or-nothing stance on latency, with adherence to SLOs considered complete success, and violations viewed as failure. Accordingly, these platforms schedule inference jobs in a work-conserving manner and select the max batch size that limits SLO violations for queued requests. However, many interactive applications present a more nuanced latency story where sub-SLO responses are not equally useful, e.g., faster responses boost conversational interactivity for chatbots (Hinton et al., 2012; yiin Chang et al., 2020) and confidence in scene perception for video analytics (Suprem et al., 2020; Bhardwaj et al., 2022).
Other platforms (Olston et al., 2017; tri, 2024; Tor, 2024) provide more flexibility by exposing tunable knobs to guide queue management, e.g., $max\_batch\_size$ and $batch\_timeout\_micros$ parameters cap batch sizes or inter-job scheduling durations. However, such knobs do little to ease the throughput-latency tension, presenting harsh tradeoffs (Figure 2): tuning for median latency improvements of 17.3–39.1% brings 1.1–3.6$\times$ reductions in average batch sizes (and proportional hits on throughput).
Platforms for serving generative models (Kwon et al., 2023; Yu et al., 2022) face similar tensions despite the less explicit focus on SLOs (since sequence lengths are hard to predict). Indeed, although such platforms use continuous batching to ensure that new requests immediately leverage idle resources as any input’s generation finishes, they prioritize throughput by running at the highest possible batch size (capped by a preset max).
2.2. Early-Exit Models
Early (or multi) exit models (Teerapittayanon et al., 2017; Xin et al., 2020) present an alternate way to address this tension by rethinking the granularity of inference. As shown in Figure 3, the key premise is that certain ‘easy’ inputs may not require the full predictive power of a model to generate an accurate result. Instead, results for such inputs may be predictable from the values at intermediate layers. In such cases, the foregone model execution can yield proportional reductions in both per-request latencies and compute footprints. Thus, the goal with early exits (EEs) is to determine, on a per-input basis, the earliest model layer at which an accurate response can be generated.
@@ 2.2. Early-Exit Models
Figure 4. EEs can lower latencies without harming throughput. Results modulate latencies from TF-Serve with original/vanilla models (Figure 2) based on optimal exiting.
Potential benefits. To understand the effect that EEs can have on the latency-throughput tension, we used off-the-shelf EE variants for the models in Figure 2: BranchyNet (Teerapittayanon et al., 2017) (CV) and DeeBERT (Xin et al., 2020) (NLP). For each model-input pair, we identified the optimal exit point defined as the earliest exit ramp that predicted the correct response for the input. We then modified the highest-throughput results in Figure 2 to account for exiting by subtracting the time saved for each exiting input, i.e., the difference in time for passing an input to the end of its optimal ramp versus passing it to the end of the model (without any ramps). These results are conservative upper bounds in that they do not reduce queuing delays or alter job scheduling. As shown in Figure 4, without changing queueing decisions, EEs can bring 35–54.7% and 17.9–26% improvements in median and 95th percentile latencies.
2.3. Challenges
Despite many EE proposals from the ML community (Zhou et al., 2020; Liu et al., 2020; Xin et al., 2020; Huang et al., 2018a; Teerapittayanon et al., 2017; Xin et al., 2021; Schwartz et al., 2020; Bae et al., 2023; Schuster et al., 2022), and their potential benefits, multiple issues complicate EE use in practice, limiting adoption.
C1: Latency and resource overheads. Although exiting can enable certain inputs to eschew downstream model computations, exit ramps impose two new overheads on serving. First, to be used, ramps must also be loaded into GPU memory which is an increasingly precious resource as models grow in size (Yu et al., 2022; Sevilla et al., 2021; Jeon et al., 2019) and inference spreads to resource-constrained settings (Padmanabhan et al., 2022; Guo et al., 2021). For instance, DeeBERT inflates memory requirements by 6.6% compared to BERT-base. Second, certain inputs may be too “hard” to accurately exit at an intermediate ramp. In these cases, serving latency and throughput mildly worsen as unsuccessful exiting decisions are made, e.g., inputs that cannot exit at any ramp slow by 22.0% and 19.5% with BranchyNet and DeeBERT.